OpenStudy (princesssleelee):

Based on the table of values below, find the slope between points where x = 1 and where x = 4. x y 1 8 3 6 4 −1 −3 negative two thirds three over two 3

OpenStudy (princesssleelee):

@igigighjkl

OpenStudy (princesssleelee):

lol im sorry it wrote out all oddly

OpenStudy (alexandervonhumboldt2):

yeah i don't understand it

OpenStudy (alexandervonhumboldt2):

can you post a screenshot?

OpenStudy (princesssleelee):

I will try

OpenStudy (princesssleelee):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

thats better thx: use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }=\frac{ 8+1}{ 1-4}=\frac{ 9 }{ -3 }=?\]